Bobby and Rick are in a 10-lap race on a one-mile track. Bobby, averaging 90mph, has completed two laps just as Rick is getting his car onto the track. What speed does Rick have to average to be even with Bobby at the end of the tenth lap?
Answer by longjonsilver(2297)

Bobby has 8 laps to travel: 8 miles going at 90mph. How long will he take to finish?

speed = distance/time
90 = 8/t
90t = 8
t = 8/90

No point working this out yet... it is a recurring decimal, so do not round it. Leave it as a fraction.

So, Rick comes onto the track. He has 10 laps to do ie 10 miles in the same time, to finish equal with Bobby.

So, speed = distance/time
speed = 10/(8/90)
speed = 900/8
speed = 112.5mph

jon.